a sample oh hydrogen gas, H2(g), is placed in a 0.500L container at 295 K. The gas pressure is 1.442bar. how many moles of H2 gas are in the container
1
Expert's answer
2020-03-26T03:03:17-0400

pV=nRT

T = 295 K

V = 0.5 L = 0.5*10-3 m3

p = 1.442 bar = 144200 Pa

R = 8.314 J/mol*K

n = pV/RT = (144200*0.5*10-3)/(8.314*295) = 0,029 mol.